A reaction has an equilibrium constant of 6.7×103 at 298K . At 727K , the equilibrium constant is 0.33.Solve for delta H the reaction
Answer #1

Van't Hoff equation

ln(K2/K1) = ( -ΔH / R ) ( 1/T2 - 1/T1 )

ΔH = -Rln(K2/K1) / ( 1/T2 - 1/T1 ) = -8.314ln(0.3/6700)(1/727-1/298)

ΔH= -41648 J/mol = -41.648 kJ/mol
Answer #2

vant Hoff equation: ln(K2/K1) (AH/R) x (1/T2 1/T1) T1 = 298 K, T2 = 737 K K1 6.7 x 103, K2 0.33 R = 8.314 J/mol·K In(0.33/6.